Hall locked around 6 pm but only two hours of beam this shift. Injection
current was particularly low (<50 mA) until shift change (8 pm). Then some
steering (around IT region) actually helped. 
Started out with 9 KHz at 80 mA on an empty target, trigger was 
elastic_paddles12-13. Before the collimator went in we had 100 Hz with 
this trigger at 80 mA (last week). However, the collimator is unshielded 
(so is generating background). Secondly the collimator center is 2.5 mm 
lower than nominal beam height. Note that -at least for the target exit 
window- we could confirm today that they are at the right beam height. 
Presumably also the cell. There is a "local bump" for both the collimator 
position and the IT position, however, these two bumps have 3 steering 
coils in common so changing one will affect the other. The collimator is 
just downstream of the LS2HV steerer.
The strategy was: cut the halo first, change vertical position at the 
collimator (that is intended to cut for background from the injection septum) 
in close steps, do a quick H scan at best V position, possibly add an angle 
to get the best part of the storage cell. Turn field on. I will write a 
short web-memo on how to proceed/what to expect. Unfortunately tonight I 
could only do the first part (see below). Also, this steering can affect 
the beam position at the Compton.
After playing with the slits I got 1 KHz at 80 mA on an empty target. 
Beyond that improvement was marginal and at the cost of serious 
lifetime reduction. 
Then tried to move the beam "1 mm down" .. and lost it immediately. 
Then we lost the drive system for the klystron and we were out for the night. 
Could not use the BLAST power supply with beam on. The above rates presumably 
relates to a beam far down at the collimator position which maybe was 
needed for injection. Slits settings were not too far away (within 2 mm) 
from was found out a couple of weeks ago.
We have taken first WCH data, only 16 channels. Coda still hangs if all 
channels are connected. Mostly because all channels ring multiple times.
There is room to play with the length of the window for the multihit TDC 
(0.5 - 32 us), the depth of the multihit buffer (1-16 hits), discriminator 
thresholds, HV, noise-chasing in general. This can also be done w/o beam
Taken flasher trigger data with no field (run#1009) and with BLAST field 
at 5000 A (run#1011)
*****Things we suffered from (will be addressed tomorrow)
_ We found there was beam in the ring, nobody called to check our HV 
  status (not the first time!!!!!). Chamber left sector tripped from 
  standby
_ We could not use the automated insertion of slits after a fill 
  (because of operator training)
_ We can move the beam "down 1 mm" but nobody really knows relatively
  to what (semantics)
_ Beam was clearly not the same of last night when we came on
regards,
tancredi
________________________________________________________________________________
Tancredi Botto,  		phone: +1-617-253-9204    fax: +1-617-253-9599
research scientist		MIT/Bates, 21 Manning Av   Middleton MA, 01949
^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^
This archive was generated by hypermail 2.1.2 : Mon Feb 24 2014 - 14:07:28 EST